OpenStudy (anonymous):

A 10 amp current passing through a wire .254m in length producing a side force of 5N a) Find the magnitude of B require to produce the force the closest equation I could find to solve part A is F=IL x B it is alright to omit the vector and use standard operation to solve this? Part B) If the wire is of copper with 0.003175 m in diameter, what would be the voltage drop and the power loss?

OpenStudy (anonymous):

Yes, you can disregard the vector operation (i.e. the cross product) and treat it instead as a multiplication.

OpenStudy (anonymous):

Thanks, that's what I been thinking but I haven't got around the question why. I suppose it's because it's in perpendicular? As for Part B, if anyone have a hint please pitch in.

OpenStudy (anonymous):

Yes it's because the B field is has to be perpendicular to the wire to produce a force that is also perpendicular to the wire. For part B, start with this:\[R=rho \frac{ L }{ A }\]where R is resistance; ρ is the electrical resistivity of copper; L is the wire length; and A is the cross sectional area of the wire.

OpenStudy (anonymous):

Well, that didn't work. Here it is:\[R = rho \frac{ L }{ A }\]

OpenStudy (anonymous):

That...made a lot of sense, thank you.

OpenStudy (anonymous):

Last bump? for R, I got 5.4218*10^-4 Am a bit confused on the voltage drop and power loss. I*R=V Concept wise, could anyone explain what is happening in this problem?

OpenStudy (anonymous):

You've got the right equation for voltage drop. Power in an electric circuit is given by:\[P=VI\]where P is power; V is voltage; and I is current.

OpenStudy (anonymous):

Note that that power equation has several forms:\[P=VI=\frac{ V ^{2} }{ R }=I ^{2}R\]

OpenStudy (anonymous):

I feel like am asking way too much question for this one particular question. Thank you for answering all of these. Just one last question, although I've been using the equation I*R=V a lot lately, modeling circuits and so forth, I can't seem to grasp the fundamental(?) of electromagnetic field and how voltage drop plays a role in this model.

OpenStudy (anonymous):

For the sake of understanding the concept and appreciating all these formula (heheh) could you please explain.

OpenStudy (anonymous):

Voltage drop is the same thing as voltage: it's the potential difference between two points. In this case, it's the potential difference between one side of the resistor and the other. In the case of a circuit with a battery, it's the potential difference between the battery's positive and negative terminals. Voltage can be written as:\[V=Ed\]where V is voltage; E is the electric field strength; and d is the distance between two points in the electric field. If we use the formula for electric field we get this:\[V=Ed=\left[ \frac{ 1 }{ 4\pi \epsilon _{O} }\frac{ q _{1}q _{2} }{ r ^{2} } \right]r=\frac{ 1 }{ 4\pi \epsilon _{O} }\frac{ q _{1}q _{2} }{ r}\]

OpenStudy (anonymous):

I think I see it now. Even though this deals with electromagnetic field(?) the property of voltage (by definition and practicality remain the same. THANKS!

OpenStudy (anonymous):

For this problem, it's actually not necessary to know anything about the electric field.